### v2

@@ Abstract
Key words. Nonsmooth problems, semismoothness, solution map, generalized derivative
MSC codes. 65K10, 65K15, 90C26.
@@ 1 Introduction
For solving problems of the form
@@
where $\theta:\mathbb{R}^{n}\to\mathbb{R}$ is a locally Lipschitz function and $U_{\rm ad}\subset\mathbb{R}^{n}$ is a closed set, the use of bundle methods is a popular approach. There exists a bunch of bundle methods in the literature, cf. [23, 19, 36], and all these methods have in common that they require an
oracle, which returns for arbitrary $x\in\mathbb{R}^{n}$ the function value $\theta(x)$ and one subgradient $g(x)\in\partial^{c}\theta(x)$, (1.2)
@@
From a limited number of bundle methods, applicable to the numerical solution of (1.1) (e.g. [23, 19, 36]), we consider in this paper the Bundle-Trust region algorithm (BT algorithm) from Schramm and Zowe [36], described also in [40]. It seems that Dempe and Bard [8] were the first who observed that the BT algorithm also works, when the oracle delivers only a so-called pseudogradient. Pseudogradients were introduced by Norkin [29, 30] and have the remarkable feature that they fulfill a certain semismoothness property. Although this property does not necessarily imply that $\theta$ is weakly semismooth, one can show that a similar convergence result holds as above: If the BT-algorithm produces a bounded sequence then at least one accumulation point is pseudo-stationary, i.e., 00 is a pseudogradient of $\theta$ at this point. By checking the convergence proof for other bundle methods [19, 23] one can show that a similar assertion holds: If the oracle delivers only a pseudogradient, one can still prove that the method works and returns pseudo-stationary points.
@@
the so-called semismooth Newton method was introduced in [11], where $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is a set-valued mapping. Here, the semismooth property was defined in terms of the limiting (Mordukhovich) coderivative and was later simplified to SC (subspace containing) derivative in [12]. The SC derivative can be considered as a generalization of the B-Jacobian (Bouligand) of single-valued mappings to multifunctions and acts as a kind of skeleton for the limiting coderivative. The semismooth Newton method has been already successfully applied to several difficult problems, see, e.g., [14, 15, 9, 10].
We observe that semismoothness plays an important role in solving different types of nonsmooth problems: optimization problems, equations and inclusions. The respective notions of semismoothness are slightly different for the different problem types, which complicates the numerical solution of problems assembled by different nonsmooth and set-valued objects. As an example consider an optimization problem of the form
@@
$\min\theta(x):=\varphi(x,\sigma(x))\quad\mbox{ subject to }x\in U_{\rm ad}.$
@@
In this paper we will provide the theoretical base for this approach. In particular, we will conduct a deep analysis of various available notions of the semismoothness property and find out some important mutual relationships. We will show that a single-valued mapping possesses some semismooth derivative on an open set $\Omega$ if and only if it is G-semismooth (semismooth in the sense of Gowda [16]) there. Further, this mapping is almost everywhere strictly differentiable and the semismooth derivative coincides with the generalized Jacobian almost everywhere in $\Omega$. Moreover, its generalized Jacobian is contained in another semismooth derivative obtained from the original one by performing some closure operation and taking the convex hull of the values. For set-valued mappings we introduce a generalization of the semismooth property (denoted $\Gamma$-ss property) and the so-called SCD-ss property. These new notions enable us, among other things, to analyze the semismoothness of solution maps to inclusions appearing in (1.3) by examination of the underlying mapping $F$. Note that, up to now, only results on the semismoothness property of single-valued solution mappings given implicitly by single-valued equations are available, cf. [37, 31, 16, 25, 21].
As already observed by Chen, Sun and Zhang in the very recent paper [5], there are some relationships between G-semismoothness and the semismooth property. We extend these results by showing that a graphically Lipschitzian mapping (i.e., a multifunction whose graph coincides with the graph of a Lipschitz single-valued mapping after a change of coordinates) is SCD-ss if and only if the transformed mapping is G-semismooth. As a byproduct we obtain that an SCD-ss mapping is almost everywhere strictly proto-differentiable and the SC derivative is almost everywhere a singleton showing the easiness of its computation there.
Finally, on the basis of the above theoretical results, we will derive for problem (1.3) two implementable algorithmic schemes, enabling us the computation of a semismooth derivative of $\theta$ via either the limiting coderivative or the SC derivatives of $F$. The respective oracles will then, in cooperation with a suitable bundle method, generate stationary points even in some situations outside the classical area of applicability of ImP. This is illustrated by an academic example. In a subsequent paper we intend to work out this approach to several instances of (1.3) having a practical relevance and to demonstrate its efficiency by numerical experiments.
@@
The paper is organized in the following way. Section 2 is divided into several subsections with the first two containing the basic notions of variational analysis and generalized differentiation. In Subsection 2.3 we describe a general principle, analogous to the construction of Clarke generalized Jacobian, how to generate from an arbitrary multifunction another multifunction having some favorable properties. In the next subsection we recall the basic properties of SCD (subspace containing derivative) mappings. In Subsection 2.5 we recall the semismooth property of set-valued mappings and its extension to SCD mappings.
The main new results are presented in Sections 3–5. Section 3 is devoted to the semismoothness property of single-valued mappings with respect to certain multifunctions. This concept is commonly used in semismooth Newton methods for solving nonsmooth equations, in particular in infinite dimensions. However, in contrast to these applications, we do not consider the case that semismoothness only holds at the solution, but on the whole domain. There is a close relation with the seminal work of Norkin [29, 30] on generalized differentiation of real-valued functions. In Section 4 we present important relationships between the G-semismooth and the SCD-ss property for graphically Lipschitzian multifunctions. Section 5 concerns the semismooth property of solution maps to inclusions. We establish the semismooth property of general solution maps and the semismooth property for continuous localizations and selections, respectively. Finally it will be shown that the analysis considerably simplifies for SCD mappings.
The following notation is employed. Given a linear subspace $L\subseteq\mathbb{R}^{n}$, $L^{\perp}$ denotes its orthogonal complement. Further, given a multifunction $F$, $\mathrm{gph}\,F:=\{(x,y)\,\big{|}\,y\in F(x)\}$ stands for its graph and $\mathrm{dom}\,F:=\{x\,\big{|}\,F(x)\neq\emptyset\}$ denotes its domain. For an element $u\in\mathbb{R}^{n}$, $\|u\|$ denotes its Euclidean norm and ${\cal B}_{\delta}(u)$ denotes the closed ball around $u$ with radius $\delta$. The closed unit ball is denoted by ${\cal B}$. In a product space we use the norm $\|(u,v)\|:=\sqrt{\|u\|^{2}+\|v\|^{2}}$. The space of all $m\times n$ matrices is denoted by $\mathbb{R}^{m\times n}$. For a matrix $A\in\mathbb{R}^{m\times n}$ , we employ the operator norm $\|A\|$ with respect to the Euclidean norm and we denote the range of $A$ by ${\rm rge\;}A$. Given a set $\Omega\subset\mathbb{R}^{s}$, we define the distance of a point $x$ to $\Omega$ by ${\rm dist}(x,\Omega):=\inf\{\|y-x\|\,\big{|}\,y\in\Omega\}$ and the indicator function is denoted by $\delta_{\Omega}$. Further we define for a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, and $x\in\Omega$ the moduli
@@ 2.1 Nonsmooth analysis
¯$\overline{\nabla}F(\bar{x})=\mathop{{\rm Lim}\,{\rm sup}}_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\{\nabla F(x)\},\ \bar{x}\in\Omega.$
@@
If $F$ is locally Lipschitz, by Rademacher’s Theorem the set $\Omega\setminus{\cal O}_{F}$ has Lebesgue measure 00. Since the Jacobians $\nabla F(x)$, $x\in{\cal O}_{F}$, are also locally bounded by the Lipschitz constant, we obtain ¯$\overline{\nabla}F(x)\subset\mathrm{conv}\,\overline{\nabla}F(x)$, $x\in\Omega$. Further, we say that $F$ is semismooth at $\bar{x}\in\Omega$ if $F$ is directionally differentiable at $\bar{x}$ and
@@
    ¯$\overline{\nabla}F(\bar{x})$ is a singleton.
@@
ad (i) $\Rightarrow$ (iii): If $F$ is strictly differentiable at $\bar{x}$ then $\nabla F$ is continuous at $\bar{x}$ relative to ${\cal O}_{F}$, cf. [35, Exercise 9.64]. Thus $\lim_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists (and equals to $\nabla F(\bar{x})$).
@@
ad (ii) $\Rightarrow$ (i): Let ¯$\overline{\nabla}F(\bar{x})=\{A\}$. For every $i=1,\ldots,m$, we conclude from [6, Theorem 2.6.6] that the $i$-th component of $F$, $i=1,\ldots m$, satisfies ¯$\bar{\nabla}F_{i}(\bar{x})=\{A_{i}\}$, where $A_{i}$ denotes the $i$-th row of $A$. Thus $\partial^{c}F_{i}(\bar{x})$ is a singleton and therefore $F_{i}$ is strictly differentiable at $\bar{x}$ by [6, Proposition 2.2.4]. Since every component $F_{i}$, $i=1,\ldots m$, is strictly differentiable at $\bar{x}$, so is $F$ as well. ∎
@@ 2.2 Variational geometry and differentiation of multifunctions
$\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})\atop t\downarrow 0}}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{{{\textstyle{(\atop x}},y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}{t\downarrow 0}}\frac{\mathrm{gph}\,F-(x,y)}{t}\big{(}=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})\big{)}.$ (2.5)
@@
By taking into account [35, Theorem 6.26], we obtain from (2.5) that
@@
In case of single-valued Lipschitz maps, strict proto-differentiability coincides with strict differentiability.
Lemma 2.5.
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $F:\Omega\to\mathbb{R}^{m}$ be Lipschitz near $\bar{x}\in\Omega$. Then $F$ is strictly proto-differentiable at $(\bar{x},F(\bar{x}))$ if and only if $F$ is strictly differentiable at $\bar{x}$.
Proof.
By Lipschitz continuity of $F$, for all convergent sequences $t_{k}\downarrow 0$, $u_{k}\to u$ and $x_{k}\to\bar{x}$ the sequence $(F(x_{k}+t_{k}u_{k})-F(x_{k}))/t_{k}$ has at least one accumulation point. Thus, strict proto-differentiability of $F$ at $\bar{x}$ is equivalent with the requirement that for such sequences $t_{k}$, $u_{k}$, $x_{k}$ the limits
$\lim_{k\to\infty}\frac{F(x_{k}+t_{k}u_{k})-F(x_{k})}{t_{k}}=\lim_{k\to\infty}\frac{F(x_{k}+t_{k}u)-F(x_{k})}{t_{k}}$
exist and are equal to $D_{*}F(\bar{x})(u)$. Hence, if $F$ is strictly proto-differentiable at $(\bar{x},F(\bar{x}))$ then $D_{*}F(\bar{x})$ is single-valued, $\mathrm{dom}\,D_{*}F(\bar{x})=\mathbb{R}^{n}$ and $\mathrm{gph}\,D_{*}F(\bar{x})$ is a subspace by Lemma 2.4. We conclude that $D_{*}F(\bar{x},\bar{y})$ is in fact a linear mapping from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ and therefore there exists an $m\times n$ matrix $A$ with $D_{*}F(\bar{x})(u)=Au$, $u\in\mathbb{R}^{n}$. Thus $F$ is strictly differentiable at $\bar{x}$. The reverse implication easily follows from the definition of strict differentiability and (2.4). ∎
@@ 2.3 Closure of mappings
It is well-known that the Clarke subdifferential $\partial^{c}\theta$ of a Lipschitz continuous function $\theta:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, is upper semicontinuous (usc) and has nonempty, convex and compact values $\partial^{c}\theta(x)$ for every $x\in\Omega$. We now discuss a possibility how to extend an arbitrary locally bounded mapping to a mapping having these properties.
@@
We denote by $\mathrm{bnd}\,\Psi(\bar{x})$ the infimum of all reals $M$ such that (2.7) holds for some neighborhood $U$. By the definition, the function $x\mapsto\mathrm{bnd}\,\Psi(x)$ is usc. Further, we denote by ${\rm cocl}\,\Psi$ the mapping defined by
@@
Example 2.6.
Let $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, be locally Lipschitz and consider the mapping $\Psi_{F}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ given by $\Psi_{F}(x):=\{\nabla F(x)\},\ x\in\mathrm{dom}\,\Psi_{F}:={\cal O}_{F}$. Then for every $x\in\Omega$ there holds $(\mathrm{cl}\,\Psi_{F})(x)=\overline{\nabla}F(x)$ and $({\rm cocl}\,\Psi_{F})(x)=\mathrm{conv}\,\overline{\nabla}F(x)$.

### v3

@@ Abstract
Key words. Nonsmooth problems, semismoothness, solution map, generalized derivative MSC codes. 65K10, 65K15, 90C26.
@@ 1 Introduction
As stated in numerous works in the literature, semismoothness plays an important role in solving various nonsmooth problems. Let us list three typical problem classes of this sort.
  1. 1.
    The first one provides the main motivation for this paper and concerns optimization problems of the form
@@
    where $\theta:\mathbb{R}^{n}\to\mathbb{R}$ is a locally Lipschitz function and $U_{\rm ad}\subset\mathbb{R}^{n}$ is a closed set. For the numerical solution of these problems, the use of bundle methods is a popular approach. There exists a bunch of bundle methods in the literature, cf. [23, 19, 37], and all these methods have in common that they require an
    oracle, which returns for arbitrary $x\in U_{\rm ad}$ the function value $\theta(x)$ and one subgradient $g(x)\in\partial^{c}\theta(x)$, (1.2)
@@
    From a limited number of bundle methods, applicable to the numerical solution of (1.1) (e.g. [23, 19, 37]), we consider in this paper the Bundle-Trust region algorithm (BT algorithm) from Schramm and Zowe [37], described also in [41]. It seems that Dempe and Bard [8] were the first who observed that the BT algorithm also works, when the oracle delivers only a so-called pseudogradient. Pseudogradients were introduced by Norkin [29, 30] and have the remarkable feature that they fulfill a certain semismoothness property. Although this property does not necessarily imply that $\theta$ is weakly semismooth, one can show that a similar convergence result holds as above: If the BT-algorithm produces a bounded sequence then at least one accumulation point is pseudo-stationary, i.e., 0 is a pseudogradient of $\theta$ at this point. By checking the convergence proof for other bundle methods [19, 23] one can show that a similar assertion holds: If the oracle delivers only a pseudogradient, one can still prove that the method works and returns pseudo-stationary points.
  2. 2.
@@
    the so-called semismooth Newton method was introduced in [11], where $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is a set-valued mapping. Here, the semismooth property was defined in terms of the limiting (Mordukhovich) coderivative and was later simplified for the important subclass of SCD (subspace containing derivatives) mappings by using the notion of SC (subspace containing) derivatives in [12]. The SC derivative can be considered as a generalization of the B-Jacobian (Bouligand differential) of single-valued mappings to multifunctions and acts as a kind of skeleton for the limiting coderivative. The semismooth Newton method has been already successfully applied to several difficult problems, see, e.g., [14, 15, 9, 10].
We observe that semismoothness plays an important role in solving different types of nonsmooth problems: optimization problems, equations and inclusions. The respective notions of semismoothness are slightly different for the different problem types and it is not clear how they can combined for the numerical solution of problems assembled by different nonsmooth and set-valued objects. As an example consider an optimization problem of the form
@@
$\min\theta(x):=\varphi(x,\sigma(x))\quad\mbox{ subject to }x\in U_{\rm ad}.$ (1.4)
@@
In this paper we will provide the theoretical base for this approach. In particular, we will conduct a deep analysis of various available notions of the semismoothness property and find out some important mutual relationships. We will show that a single-valued mapping possesses some semismooth derivative on an open set $\Omega$ if and only if it is G-semismooth (semismooth in the sense of Gowda [16]) there. Further, this mapping is almost everywhere strictly differentiable and the semismooth derivative coincides with the generalized Jacobian almost everywhere in $\Omega$. Moreover, its generalized Jacobian is contained in another semismooth derivative obtained from the original one by performing some closure operation and taking the convex hull of the values. For set-valued mappings we introduce a generalization of the semismooth property (denoted $\Gamma$-$s$ property) and the so-called SCD-$s$ property. These new notions enable us, among other things, to analyze the semismoothness of solution maps to inclusions appearing in (1.3) by examination of the underlying mapping $F$. Note that, up to now, only results on the semismoothness property of single-valued solution mappings given implicitly by single-valued equations are available, cf. [38, 31, 16, 25, 21].
As already observed by Chen, Sun and Zhang in the very recent paper [5], there are some relationships between G-semismoothness and the semismooth property. We extend these results by showing that a graphically Lipschitzian mapping (i.e., a multifunction whose graph coincides with the graph of a Lipschitz single-valued mapping after a change of coordinates) is SCD-$s$ if and only if the transformed mapping is G-semismooth. As a byproduct we obtain that an SCD-$s$ mapping is almost everywhere strictly proto-differentiable and the SC derivative is almost everywhere a singleton showing the easiness of its computation there.
Finally, on the basis of the above theoretical results, we will derive for problem (1.3) two implementable algorithmic schemes, enabling us the computation of a semismooth derivative of $\theta$ given by (1.4) via either the limiting coderivative or the SC derivatives of $F$. The respective oracles will then, in cooperation with a suitable bundle method, generate stationary points even in some situations outside the classical area of applicability of ImP. This is illustrated by an academic example. In a subsequent paper we intend to work out this approach to several instances of (1.3) having a practical relevance and to demonstrate its efficiency by numerical experiments.
@@
The paper is organized in the following way. Section 2 is divided into several subsections with the first two containing the basic notions of variational analysis and generalized differentiation. In Subsection 2.3 we describe a general principle, analogous to the construction of Clarke generalized Jacobian, how to generate from an arbitrary multifunction another multifunction having some favorable properties. In the next subsection we recall the basic properties of SCD mappings. In Subsection 2.5 we recall the semismooth property of set-valued mappings and its extension to SCD mappings.
The main new results are presented in Sections 3–5. Section 3 is devoted to the semismoothness property of single-valued mappings with respect to certain multifunctions. This concept is commonly used in semismooth Newton methods for solving nonsmooth equations, in particular in infinite dimensions. However, in contrast to these applications, we do not consider the case that semismoothness only holds at the solution, but on the whole domain. There is a close relation with the seminal work of Norkin [29, 30] on generalized differentiation of real-valued functions. In Section 4 we present important relationships between the G-semismooth and the SCD-$s$ property for graphically Lipschitzian multifunctions. Section 5 concerns the semismooth property of solution maps to inclusions. We establish the semismooth property of general solution maps and the semismooth property for continuous localizations and selections, respectively. Finally it will be shown that the analysis considerably simplifies for SCD mappings.
[…]
@@ 2.1 Nonsmooth analysis
$\overline{\nabla}F(\bar{x})=\mathop{{\rm Lim}\,{\rm sup}}_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\{\nabla F(x)\},\ \bar{x}\in\Omega.$
@@
If $F$ is locally Lipschitz, by Rademacher’s Theorem the set $\Omega\setminus{\cal O}_{F}$ has Lebesgue measure 0. Since the Jacobians $\nabla F(x)$, $x\in{\cal O}_{F}$, are also locally bounded by the Lipschitz constant, we obtain $\emptyset\not=\overline{\nabla}F(x)\subset\mathrm{conv}\,\overline{\nabla}F(x)$, $x\in\Omega$. Further, we say that $F$ is semismooth at $\bar{x}\in\Omega$ if $F$ is directionally differentiable at $\bar{x}$ and
@@
    $\overline{\nabla}F(\bar{x})$ is a singleton.
@@
ad (i) $\Rightarrow$ (iii): If $F$ is strictly differentiable at $\bar{x}$ then $\nabla F$ is continuous at $\bar{x}$ relative to ${\cal O}_{F}$, cf. [36, Exercise 9.64]. Thus $\lim_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists (and equals to $\nabla F(\bar{x})$).
@@
ad (ii) $\Rightarrow$ (i): Let $\overline{\nabla}F(\bar{x})=\{A\}$. For every $i=1,\ldots,m$, we conclude from [6, Theorem 2.6.6] that the $i$-th component of $F$, $i=1,\ldots m$, satisfies $\bar{\nabla}F_{i}(\bar{x})=\{A_{i}\}$, where $A_{i}$ denotes the $i$-th row of $A$. Thus $\partial^{c}F_{i}(\bar{x})$ is a singleton and therefore $F_{i}$ is strictly differentiable at $\bar{x}$ by [6, Proposition 2.2.4]. Since every component $F_{i}$, $i=1,\ldots m$, is strictly differentiable at $\bar{x}$, so is $F$ as well. ∎
@@ 2.2 Variational geometry and differentiation of multifunctions
$\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})\atop t\downarrow 0}}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})\atop t\downarrow 0}}}\frac{\mathrm{gph}\,F-(x,y)}{t}\big(=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})\big).$ (2.6)
@@
By taking into account [36, Theorem 6.26], we obtain from (2.6) that
@@ 2.3 Closure of mappings
It is well-known that the Clarke subdifferential $\partial^{c}\theta$ of a Lipschitz continuous function $\theta:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, is upper semicontinuous (usc) and has nonempty, convex and compact values $\partial^{c}\theta(x)$ for every $x\in\Omega$. We now discuss a possibility of how to extend an arbitrary locally bounded mapping to a mapping having these properties.
@@
We denote by $\mathrm{bnd}\,\Psi(\bar{x})$ the infimum of all nonnegative reals $M$ such that (2.8) holds for some neighborhood $U$ with the convention $\mathrm{bnd}\,\Psi(\bar{x}):=\infty$ if $\Psi$ is not locally bounded at $\bar{x}$. We claim that the function $\mathrm{bnd}\,\Psi:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ is usc, i.e., $\limsup_{x\to\bar{x}}\mathrm{bnd}\,\Psi(x)\leq\mathrm{bnd}\,\Psi(\bar{x})$, $\bar{x}\in\mathbb{R}^{n}$. Indeed, fix any $\bar{x}\in\mathbb{R}^{n}$. If $\mathrm{bnd}\,\Psi(\bar{x})=\infty$ there is nothing to prove. If $\mathrm{bnd}\,\Psi(\bar{x})<\infty$ for every $\epsilon>0$ we can find an open neighborhood $U$ of $\bar{x}$ such that (2.8) holds with $M=\mathrm{bnd}\,\Psi(\bar{x})+\epsilon$. Then it immediately follows that $\mathrm{bnd}\,\Psi(x)\leq\mathrm{bnd}\,\Psi(\bar{x})+\epsilon$, $x\in U$, verifying upper semicontinuity of $\mathrm{bnd}\,\Psi$ at $\bar{x}$ because $\epsilon>0$ can be taken arbitrarily small.
Further, we denote by ${\rm cocl}\,\Psi$ the mapping defined by